# CLHA: A Simple yet Effective Contrastive Learning Framework for Human Alignment

### v1

@@ Abstract
| Feiteng Fang, Liang Zhu, Min Yang,††thanks: Corresponding author. Xi Feng, Jinchang Hou, |
| Qixuan Zhao, Chengming Li, Xiping Hu, Ruifeng Xu |
| University of Science and Technology of China Southern University of Science and Technology |
| Shenzhen Institute of Advanced Technology, Chinese Academy of Sciences |
| Shenzhen MSU-BIT University Harbin Institute of Technology (Shenzhen) |
@@ 3.1. Datasets
We conduct extensive experiments on the Human Preference Data about Helpfulness and Harmlessness (i.e., HH-RLHF), introduced by Bai et al. (2022). HH-RLHF has four themed sub-sets, including Harmless, Helpful, Helpful and Helpful. Each sub-set is neatly organized into distinct train/test splits. The statistics of the datasets are provided in Table 1. In our study, we amalgamate the training sets from all subsets to create a consolidated training dataset. For clarity in subsequent descriptions, we denote this original dataset as HH-RLHF, where subscript denotes the length of the generation rankings. Similar to previous study Song et al. (2023), we partition the testing set into two segments, allocating them for validation and testing purposes, respectively. Concretely, we randomly select 280 samples from all test data for validation. To rigorously evaluate the sequence data, we employed an augmented dataset following Song et al. (2023). This augmentation, HH-RLHF, involves appending a response generated by ChatGPT (gpt-3.5-turbo) to each prompt in the HH-RLHF dataset.
Subset |
#Training dataset | #Test dataset |
Harmless |
42,537 | 2,312 |
Helpful |
43,835 | 2,354 |
Helpful |
22,007 | 1,137 |
Helpful |
52,421 | 2,749 |
| Method | Harmless | Helpful | Helpful | Helpful | Total | |||||
@@
| RLHF | 14.54 | 55.05 | 19.86 | 42.16 | 28.04 | 53.40 | 25.11 | 47.73 | 21.19 | 48.93 |
| CoH | 13.34 | 45.47 | 23.17 | 39.03 | 33.84 | 52.63 | 29.79 | 46.57 | 24.06 | 45.00 |
| RRHF | 13.49 | 53.98 | 18.76 | 48.23 | 30.68 | 56.44 | 24.95 | 52.51 | 20.91 | 52.25 |
| PRO | 12.05 | 62.96 | 20.83 | 48.51 | 28.75 | 59.02 | 27.17 | 53.28 | 21.54 | 55.35 |
| RLHF | 13.63 | 61.97 | 20.12 | 55.29 | 28.89 | 59.78 | 24.65 | 58.26 | 20.99 | 58.65 |
| CoH | 13.44 | 56.87 | 21.89 | 51.52 | 34.04 | 59.51 | 28.24 | 56.35 | 23.26 | 55.58 |
| RRHF | 13.02 | 64.63 | 18.95 | 61.38 | 31.37 | 63.26 | 24.75 | 63.28 | 20.86 | 63.12 |
| PRO | 15.53 | 73.08 | 22.30 | 64.78 | 29.35 | 66.66 | 27.49 | 66.95 | 23.07 | 67.97 |
| CLHA | 15.09 | 72.88 | 22.42 | 65.13 | 30.13 | 67.45 | 27.49 | 67.49 | 23.01 | 68.30 |
@@ 4.1. Main Results
Table 2 presents the BLEU and reward model scores of our method and the compared baselines over two datasets. On the HH-RLHF dataset, CLHA outperforms all baseline methods in terms of open-source reward model scores. Specifically, CLHA achieves a 2.37% gain over the best-performing baseline, PRO. Additionally, CLHA manages to outshine other methods across the four subsets with performance improvements ranging from 1% to 4%. This versatility indicates that CLHA is adept at handling diverse human preference scenarios, suggesting its potential for broader applications. Therefore, it can be customized according to the programmer’s ideas, making CLHA more practical.
When it comes to the BLEU metric, which assesses the fluency and relevance of the generated responses, CLHA delivers consistent performances. Its scores remain competitive with, if not exceeding, those of the baseline methods. This is crucial, as it suggests that while CLHA is effective in generating aligned content with human preferences, it maintains the quality of responses, a trait inherent from the fine-tuning phase.
Furthermore, upon the introduction of augmented data from ChatGPT-3.5 into the HH-RLHF dataset, we have observed a noticeable improvement in the performance of CLHA, referred to as CLHA. The BLEU and reward scores for CLHA have reached up to 23.01% and 68.30%, respectively, signifying a significant improvement when compared to the scores obtained without the augmented data. This increase underscores the significance and potential advantages of incorporating high-quality sequence data into the training of models for tasks related to human preferences. These findings are in alignment with those presented in the research conducted by Wang et al. (2023), which emphasizes the value of diverse training data in enhancing model performance.
@@ 4.2. Human Evaluation
For the assessment of models in terms of human alignment with preferences, direct human evaluation is widely acknowledged as the most robust method. To this end, we randomly select 300 instances from each test subset within HH-RLHF and present the manually annotated results, as depicted in Figure 3. Both CLHA and PRO models are trained under identical settings using the raw HH-RLHF dataset. This dataset, characterized by a sequence length of 2, results in the derived models CLHA and PRO. The results reveal a clear superiority of responses generated by CLHA over those produced by PRO, providing robust evidence for the effectiveness of our proposed approach.
@@ 4.3. Ablation Study
To discern the individual contributions of various components to our contrastive learning framework, we conduct ablation studies on the raw HH-RLHF dataset (HH-RLHF), as detailed in Table 3. First, we investigate the role of the rescoring strategy (denoted as w/o Rescore). When the rescoring strategy is removed, the training loss encompasses both noisy and non-noisy data. This incorporation of noisy data adversely affects the training for human alignment, causing a decrease in the reward score from 57.72% to 55.48%. Next, we assess the significance of the pair-wise contrastive loss by excluding it (denoted as w/o ). Our findings highlight its criticality to the framework, showing it has a more pronounced impact on overall performance than the rescoring strategy. Lastly, based on Table 3, the margin term enhances the ability of the pair-wise contrastive loss to regulate sample generation probability. Excluding the margin term (denoted w/o ) results in a decline in the effectiveness of the method by 1.03%.
@@ 4.4. Case Study
In this section, we provide a detailed examination of two cases taken from Harmless and Helpful to evaluate the performance of CLHA. Refer to Table 4 for the original user queries and the respective responses from CLHA and CLHA. For a more comprehensive analysis, we also include the results after omitting the contrastive learning loss, denoted as CLHA w/o .
In the first case, the absence of the contrastive learning loss in the model leads to difficulties in accurately understanding the query. In comparison, CLHA employs a strategy to avoid harmful queries, aiming to reduce the likelihood of generating unsafe responses. Meanwhile, CLHA offers a response that directly addresses the unethical intent behind the query “steal a wallet from a lady”, stressing the legal and ethical implications of such actions. The answer is coherent and ensured the dissemination of accurate information.
For the second case, all three models appears to produce coherent answers. However, a deeper examination reveals differences in their effectiveness. The model without the contrastive learning loss fails to recognize the important concept of “cats having a sense of territory”. As a result, its response, while coherent, lacks key information. On the other hand, both CLHA models identifies the main point of the query. However, the response from CLHA is longer than necessary. CLHA, in contrast, provides a concise and accurate response, making it the most informative of the three.
@@ Acknowledgements
[…]

### v2

@@ Abstract
Feiteng Fang1,3, Liang Zhu2,3, Min Yang3†,††thanks: †Corresponding author. Xi Feng1,3, Jinchang Hou1,3, |
Qixuan Zhao1,3, Chengming Li4, Xiping Hu4, Ruifeng Xu5 |
1University of Science and Technology of China 2Southern University of Science and Technology |
3Shenzhen Institute of Advanced Technology, Chinese Academy of Sciences |
4Shenzhen MSU-BIT University 5Harbin Institute of Technology (Shenzhen) |
@@ 3.1. Datasets
We conduct extensive experiments on the Human Preference Data about Helpfulness and Harmlessness (i.e., HH-RLHF), introduced by Bai et al. (2022). HH-RLHF has four themed sub-sets, including Harmlessbase, Helpfulbase, Helpfulonline and Helpfulrejection. Each sub-set is neatly organized into distinct train/test splits. The statistics of the datasets are provided in Table 1. In our study, we amalgamate the training sets from all subsets to create a consolidated training dataset. For clarity in subsequent descriptions, we denote this original dataset as HH-RLHF2, where subscript denotes the length of the generation rankings. Similar to previous study Song et al. (2023), we partition the testing set into two segments, allocating them for validation and testing purposes, respectively. Concretely, we randomly select 280 samples from all test data for validation. To rigorously evaluate the sequence data, we employed an augmented dataset following Song et al. (2023). This augmentation, HH-RLHF3, involves appending a response generated by ChatGPT (gpt-3.5-turbo) to each prompt in the HH-RLHF dataset.
Subset | #Training dataset | #Test dataset |
Harmlessbase | 42,537 | 2,312 |
Helpfulbase | 43,835 | 2,354 |
Helpfulonline | 22,007 | 1,137 |
Helpfulrejection | 52,421 | 2,749 |
Method | Harmlessbase | Helpfulbase | Helpfulonline | Helpfulrejection | Total | |||||
@@
RLHF2 | 14.54 | 55.05 | 19.86 | 42.16 | 28.04 | 53.40 | 25.11 | 47.73 | 21.19 | 48.93 |
CoH2 | 13.34 | 45.47 | 23.17 | 39.03 | 33.84 | 52.63 | 29.79 | 46.57 | 24.06 | 45.00 |
RRHF2 | 13.49 | 53.98 | 18.76 | 48.23 | 30.68 | 56.44 | 24.95 | 52.51 | 20.91 | 52.25 |
PRO2 | 12.05 | 62.96 | 20.83 | 48.51 | 28.75 | 59.02 | 27.17 | 53.28 | 21.54 | 55.35 |
CLHA2 | 13.63 | 63.14 | 20.36 | 52.36 | 28.94 | 61.08 | 27.11 | 56.37 | 21.85 | 57.72 |
RLHF3 | 13.63 | 61.97 | 20.12 | 55.29 | 28.89 | 59.78 | 24.65 | 58.26 | 20.99 | 58.65 |
CoH3 | 13.44 | 56.87 | 21.89 | 51.52 | 34.04 | 59.51 | 28.24 | 56.35 | 23.26 | 55.58 |
RRHF3 | 13.02 | 64.63 | 18.95 | 61.38 | 31.37 | 63.26 | 24.75 | 63.28 | 20.86 | 63.12 |
PRO3 | 15.53 | 73.08 | 22.30 | 64.78 | 29.35 | 66.66 | 27.49 | 66.95 | 23.07 | 67.97 |
CLHA3 | 15.09 | 72.88 | 22.42 | 65.13 | 30.13 | 67.45 | 27.49 | 67.49 | 23.01 | 68.30 |
@@ 4.1. Main Results
Table 2 presents the BLEU and reward model scores of our method and the compared baselines over two datasets. On the HH-RLHF2 dataset, CLHA outperforms all baseline methods in terms of open-source reward model scores. Specifically, CLHA achieves a 2.37% gain over the best-performing baseline, PRO. Additionally, CLHA2 manages to outshine other methods across the four subsets with performance improvements ranging from 1% to 4%. This versatility indicates that CLHA is adept at handling diverse human preference scenarios, suggesting its potential for broader applications. Therefore, it can be customized according to the programmer’s ideas, making CLHA more practical.
When it comes to the BLEU metric, which assesses the fluency and relevance of the generated responses, CLHA2 delivers consistent performances. Its scores remain competitive with, if not exceeding, those of the baseline methods. This is crucial, as it suggests that while CLHA2 is effective in generating aligned content with human preferences, it maintains the quality of responses, a trait inherent from the fine-tuning phase.
Furthermore, upon the introduction of augmented data from ChatGPT-3.5 into the HH-RLHF3 dataset, we have observed a noticeable improvement in the performance of CLHA, referred to as CLHA3. The BLEU and reward scores for CLHA3 have reached up to 23.01% and 68.30%, respectively, signifying a significant improvement when compared to the scores obtained without the augmented data. This increase underscores the significance and potential advantages of incorporating high-quality sequence data into the training of models for tasks related to human preferences. These findings are in alignment with those presented in the research conducted by Wang et al. (2023), which emphasizes the value of diverse training data in enhancing model performance.
@@ 4.2. Human Evaluation
For the assessment of models in terms of human alignment with preferences, direct human evaluation is widely acknowledged as the most robust method. To this end, we randomly select 300 instances from each test subset within HH-RLHF2 and present the manually annotated results, as depicted in Figure 3. Both CLHA and PRO models are trained under identical settings using the raw HH-RLHF2 dataset. This dataset, characterized by a sequence length of 2, results in the derived models CLHA2 and PRO2. The results reveal a clear superiority of responses generated by CLHA2 over those produced by PRO2, providing robust evidence for the effectiveness of our proposed approach.
@@ 4.3. Ablation Study
To discern the individual contributions of various components to our contrastive learning framework, we conduct ablation studies on the raw HH-RLHF dataset (HH-RLHF2), as detailed in Table 3. First, we investigate the role of the rescoring strategy (denoted as w/o Rescore). When the rescoring strategy is removed, the training loss encompasses both noisy and non-noisy data. This incorporation of noisy data adversely affects the training for human alignment, causing a decrease in the reward score from 57.72% to 55.48%. Next, we assess the significance of the pair-wise contrastive loss by excluding it (denoted as w/o ). Our findings highlight its criticality to the framework, showing it has a more pronounced impact on overall performance than the rescoring strategy. Lastly, based on Table 3, the margin term enhances the ability of the pair-wise contrastive loss to regulate sample generation probability. Excluding the margin term (denoted w/o ) results in a decline in the effectiveness of the method by 1.03%.
@@ 4.4. Case Study
In this section, we provide a detailed examination of two cases taken from Harmlessbase and Helpfulbase to evaluate the performance of CLHA. Refer to Table 4 for the original user queries and the respective responses from CLHA2 and CLHA3. For a more comprehensive analysis, we also include the results after omitting the contrastive learning loss, denoted as CLHA2 w/o .
In the first case, the absence of the contrastive learning loss in the model leads to difficulties in accurately understanding the query. In comparison, CLHA2 employs a strategy to avoid harmful queries, aiming to reduce the likelihood of generating unsafe responses. Meanwhile, CLHA3 offers a response that directly addresses the unethical intent behind the query “steal a wallet from a lady”, stressing the legal and ethical implications of such actions. The answer is coherent and ensured the dissemination of accurate information.
For the second case, all three models appears to produce coherent answers. However, a deeper examination reveals differences in their effectiveness. The model without the contrastive learning loss fails to recognize the important concept of “cats having a sense of territory”. As a result, its response, while coherent, lacks key information. On the other hand, both CLHA models identifies the main point of the query. However, the response from CLHA2 is longer than necessary. CLHA3, in contrast, provides a concise and accurate response, making it the most informative of the three.
@@ Acknowledgements
- Bai et al. (2022) Yuntao Bai, Andy Jones, Kamal Ndousse, Amanda Askell, Anna Chen, Nova DasSarma, Dawn Drain, Stanislav Fort, Deep Ganguli, Tom Henighan, et al. 2022. Training a helpful and harmless assistant with reinforcement learning from human feedback. arXiv preprint arXiv:2204.05862.